Find the distance d(P1,P2)between the points P1 and P2.
P1=(5,-5); P2=(4,4)
Please show all of the steps..Thank you

Answer by amarjeeth123(574)

The distance between the points P1(5,-5) and P2(4,4) is d=√(5-4)^2+(-5-4)^2=√1+81=√82 units